OpenStudy (anonymous):

A certain particle has a lifetime of 1x10^-7s when measured at rest. How far does it go before decaying, its speed is 0.99c when it is created?

OpenStudy (anonymous):

this question is on time dialation, please anyone help.

OpenStudy (nottim):

Ha. I almost thought we could use d=vt here.

OpenStudy (nottim):

d=? v=0.99c t=1x10^-7s

OpenStudy (nottim):

obviously more complex than that though, eh?

OpenStudy (anonymous):

lifetime at rest is give, we have to find lifetime in motion

OpenStudy (nottim):

aww bugger. I really wanted this to be that simple.

OpenStudy (anonymous):

answer is not coming correct by your expression?

OpenStudy (anonymous):

its a time dialation problem related to relativity theory

OpenStudy (nottim):

Aww.. How does that make this more complex? Is there like some super tricky thhing going on?

OpenStudy (nottim):

heyy?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

we use this eaqution to find t\[t=t_{o} / \sqrt(1-v^2/c^2)\]

OpenStudy (anonymous):

then we use d=vt

OpenStudy (nottim):

Please be plug and play?

OpenStudy (anonymous):

what

OpenStudy (nottim):

just plug in variableS?

OpenStudy (nottim):

@salini can you help?

OpenStudy (nottim):

@Mani Jha or @Espex please help? MY god I can not do this...

OpenStudy (nottim):

yeah..I can't help you. I'm sorry.

OpenStudy (anonymous):

life time at rest = 10^(-7) s life time at speed 0.99c = 10^(-7)/sqrt(1-(0.99*0.99)) = 7.0888 x 10^(-7) s so, before decaying completely it travels 0.99c x 7.0888 x 10^(-7) m = 210.5377 m